# appendixtocaptAppendix

### v3

@@ Abstract
The hippocampus enables spatial navigation through place cell populations forming cognitive maps. We propose proximity-preserving neural embeddings to encode multi-scale random walk transitions, where the inner product $\langle h(x,t),h(y,t)\rangle=q(y|x,t)$ represents normalized transition probabilities, with $h(x,t)$ as the embedding at location $x$ and $q(y|x,t)$ as the transition probability at scale $\sqrt{t}$. This scale hierarchy mirrors hippocampal dorsoventral organization. The embeddings $h(x,t)$ reduce pairwise spatial proximity into an environmental map, with Euclidean distances preserving proximity information. We use gradient ascent on $q(y|x,t)$ for straight-forward path planning, employing adaptive scale selection for trap-free, smooth trajectories, equivalent to minimizing embedding space distances. Matrix squaring ($P_{2t}=P_{t}^{2}$) efficiently builds global transitions from local ones ($P_{1}$), enabling preplay-like shortcut prediction. Experiments demonstrate localized place fields, multi-scale tuning, adaptability, and remapping, achieving robust navigation in complex environments. Our biologically plausible framework, extensible to theta-phase precession, unifies spatial and temporal coding for scalable navigation.
@@ 1 Introduction
Hippocampal place cells are central to spatial cognition, firing selectively at specific locations during navigation [45, 46]. We reconceptualize place cells, shifting from individual place fields to the whole population as proximity-preserving neural embeddings that collectively encode multi-scale spatial relationships, forming cognitive maps [59]. This population-level view, akin to viewing a tree rather than individual leaves, captures the hippocampus’s dynamic navigation capabilities.
Place cells adapt firing fields to environmental changes [42], scale along the dorsoventral axis [31], and predict paths via preplay before exploration [13], enabling robust navigation across diverse settings and resilience to deformations [14, 49, 41]. Our framework models these traits using embeddings $h(x,t)$, naturally arising from spectral analysis of multi-time random walk transition probabilities $q(y|x,t)$ and non-negative matrix factorization for biological plausibility [31]. The inner product $\langle h(x,t),h(y,t)\rangle=q(y|x,t)$ approximates the normalized transition probability of a symmetric random walk with reflecting boundaries, where $q(x|x,t)=1$, $\|h(x,t)\|=1$, and multi-scale proximity is defined by transitions over time duration $t$, with $\sqrt{t}$ as the spatial scale. Despite being purely random, the walk’s transitions encode geodesic distance for short $t$ and topological connectivity for long $t$, capturing environmental structure. Efficient matrix squaring ($P_{2t}=P_{t}^{2}$) computes global transitions from local ones ($P_{1}$), enabling preplay-like shortcut detection [44, 14]. Non-negative constraints yield localized place fields, tiling the environment.
Navigation uses gradient ascent on $q(y|x,t)$ with adaptive scale selection, minimizing the Euclidean distance $\|h(x,t)-h(y,t)\|^{2}$ for trap-free, smooth trajectories [38, 41, 59]. We term this straight-forward path planning, as it seeks to minimize the Euclidean distance in the embedding space, simplifying complex physical navigation.
Our contributions include: (1) reconceptualizing place cells as proximity-preserving embeddings; (2) modeling multi-scale random walk transitions and derive the embeddings by spectral analysis and nonnegative matrix decomposition; (3) using matrix squaring to build global proximity relationships from local ones; (4) developing straight-forward path planning that seeks to minimize Euclidean distance in embedding space; and (5) demonstrating biological plausibility via preplay, remapping, and theta-phase precession. Bridging connectionist models [3, 62] and cognitive map theories [46, 41], our framework unifies spatial and temporal coding, extensible to theta-phase precession [47, 55].
@@ 2 Method
2.1 Multi-Time Random Walk Transition Kernel
The foundation of our approach is a symmetric random walk on a discrete lattice over the environment (e.g., a $40\times 40$ lattice), with a subset of lattice points belonging to the obstacles. This random walk serves as a mapping policy rather than a goal-reaching policy. Remarkably, this purely random exploration policy leads to optimal path planning without explicit policy optimization, because transition kernel over time contains information about geodesic distance for short time and topological connectivity for long time.
@@ 2.1 Multi-Time Random Walk Transition Kernel
  • For each unobstructed neighbor $y$ of $x$: $p(y|x,t=1)=p_{\text{move}}$ (e.g., 1/4 in the case of 4 nearest neighbors)
  • For self-transition: $p(x|x,t=1)=1-|N(x)|\cdot p_{\text{move}}$
@@
where $|N(x)|$ is the number of unobstructed neighbors of $x$. A critical property of this formulation is that the transition probabilities are symmetric, meaning $p(y|x,t=1)=p(x|y,t=1)$ for all locations $x$ and $y$. This symmetry ensures that the random walk process preserves the bidirectional nature of spatial relationships, which is essential for creating a well-defined distance metric.
@@
The above random walk defines $t$-step symmetric transition kernel $p(y|x,t)$ over time $t$, which measures the spatial proximity or adjacency between $x$ and $y$ at a spatial scale or resolution reflected by $t$. The above formulation can be applied to random walk on general graphs.
2.2 Heat Equation and Geodesic Distance
@@
Our discrete random walk model establishes a connection to the continuous heat equation. For small spatial discretization $dx$ and temporal discretization $dt$ with $dx=\sqrt{dt}$, our discrete random walk converges to the heat equation with reflecting boundary conditions as $dx\rightarrow 0$ [48, 24]:
@@
$\frac{\partial p(y|x,t)}{\partial t}=\alpha\nabla^{2}p(y|x,t)\quad\text{on}\quad\Omega\setminus\Omega_{\text{obstacles}}$ (1)
$\frac{\partial p(y|x,t)}{\partial n}=0\quad\text{on}\quad\partial\Omega_{\text{obstacles}}$ (2)
@@
where $\Omega$ is the whole region, $\Omega_{\text{obstacles}}$ is the region of obstacles, and $\partial\Omega_{\text{obstacles}}$ is its boundary. $\alpha$ is the diffusion coefficient (different one-step transition $p(y|x,t=1)$ in discrete case leads to different $\alpha$, e.g., $\alpha=1/4$ for 4-nearest neighbor random walk).
@@
A fundamental result from heat diffusion theory relates the short-time behavior of the heat kernel to geodesic distance. For small values of $t$, the heat kernel has the asymptotic form [60, 28]:
$p(y|x,t)\approx\frac{1}{4\pi\alpha t}\exp\left(-\frac{d_{g}^{2}(x,y)}{4\alpha t}\right)$ (3)
@@
where $d_{g}(x,y)$ is the geodesic distance. For open domains, $d_{g}(x,y)$ becomes Euclidean distance, and $p(y|x,t)\sim{\cal N}(x,2\alpha t)$, a Gaussian distribution with variance $2\alpha t$ or standard deviation $\sqrt{2\alpha t}$, demonstrating that $\sqrt{t}$ correspond to the spatial scale. We take square root of $t$ to emphasize this scaling relationship between time and space.
@@
This result demonstrates that $d_{t}(x,y)=-t\log q(y|x,t)$ approximates the squared geodesic distance for small values of $t$, providing a fundamental connection between random walks and geodesic distances in complex environments. As $t$ increases, the transition probability incorporates additional information about path multiplicity and global connectivity, creating a multi-scale representation of spatial relationships. In particular, for large $t$, the transition probability is dominated by eigenvalues that are close to 1, whose eigenvectors depend on topological connectedness rather than local geometry.
@@
Appendix B provides detailed background on heat equation. Appendix D explains topological properties for large $t$.
2.3 Place Cells as Position Embeddings
@@
We formalize place cell population as vector embedding $h(x,t)\in\mathbb{R}^{n}$, where $n$ is the number of place cells (e.g., $n=500$). The inner product between embeddings approximates the normalized transition probability kernel:
$\langle h(x,t),h(y,t)\rangle\approx q(y|x,t)$ (4)
@@
where $q(y|x,t)={p(y|x,t)}/{\sqrt{p(x|x,t)\cdot p(y|y,t)}}$ is the normalized transition probability, so that $q(x|x,t)=1$ and $\|h(x,t)\|=1$ for all $x$ due to normalization. For each cell $i$, $h_{i}(x,t)\geq 0$ for biological plausibility, and $h_{i}(x,t)$ is the response map or profile of cell $i$.
@@
The above formulation can be derived through spectral decomposition of the transition matrix. Since the one-step transition matrix $P_{1}$ is symmetric by construction, its powers $P_{t}=P_{1}^{t}$ admit an eigendecomposition [39]: $P_{t}=Q\Lambda^{t}Q^{T},$ where $Q$ is orthogonal ($Q^{T}Q=I$) and $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{n})$ contains eigenvalues $0\leq\lambda_{i}\leq 1$ (assuming the random walk is irreducible and aperiodic).
@@
From this spectral decomposition, we can construct a natural embedding by defining: $H_{i}(x,t)=\lambda_{i}^{t/2}Q_{i}(x)$ where $Q_{i}$ is the $i$-th column of $Q$. This yields an exact representation of the transition probability through inner products:
$p(y|x,t)=\sum_{i}H_{i}(x,t)H_{i}(y,t)=\langle H(x,t),H(y,t)\rangle$ (5)
@@
where $H_{i}(x,t)$ is the $i$-th element of the vector $H(x,t)$. For the normalized transition probability $q(y|x,t)$, we define normalized embeddings:
$h_{\text{spec}}(x,t)=\frac{H(x,t)}{\|H(x,t)\|}=\frac{H(x,t)}{\sqrt{p(x|x,t)}}$ (6)
@@
$\langle h_{\text{spec}}(x,t),h_{\text{spec}}(y,t)\rangle=\frac{p(y|x,t)}{\sqrt{p(x|x,t)\cdot p(y|y,t)}}=q(y|x,t)$ (7)
$p(x|x,t)$ is constant in the open field and is smooth in general, so $q$ is essentially a scaled version of $p$.
However, these spectral embeddings may contain negative components, conflicting with the biological constraint that neural firing rates must be non-negative. Horn’s theorem [27, 5], which is built upon the above spectral decomposition, guarantees the existence of a non-negative $h(x,t)$, such that $\langle h(x,t),h(y,t)\rangle=q(y|x,t)$ for non-negative matrix factorization.
This formulation represents a fundamental shift from modeling individual place cells to modeling the place cell population as distributed position embedding. The embedding vector $h(x,t)$ represents the firing rates of all place cells at location $x$ for spatial scale $\sqrt{t}$, capturing the idea that it is the pattern across the population—not the activity of any single cell—that encodes location.
The time parameter $\sqrt{t}$ serves as a fundamental unit of spatial resolution or scale, with larger values producing broader, more diffuse representations, and smaller values producing more localized representations. This naturally mirrors the variation in place field sizes observed along the dorsoventral axis of the hippocampus [31, 58].
Probability-geometry duality. Geometrically, $\langle h(x,t),h(y,t)\rangle$ is the cosine of the angle between the normalized vectors $h(x,t)$ and $h(y,t)$. Moreover, we have
$\frac{1}{2}\|h(x,t)-h(y,t)\|^{2}=1-2\langle h(x,t),h(y,t)\rangle=1-q(y|x,t)$ (8)
That is, the angle or the Euclidean distance between $h(x,t)$ and $h(y,t)$ encodes proximity or adjacency between $x$ and $y$. The spectral analysis or matrix decomposition naturally connects probability to geometry.
$(h(x,t),\forall x)$ is a 2D manifold in the high-dimensional embedding space. This 2D manifold is an embedding of the 2D physical space. Because of the flexibility endowed by the high-dimensional embedding space, the path between $x$ and $y$ on the manifold is essentially “straight” even though the physical path is far from being straight.
$(h(x,t),\forall x)$ can be interpreted as a map of the environment because it preserves proximity or adjacency information of the environment.
See Appendix A for more details on spectral analysis. Appendix C provides analytical results for open field, where $q(y|x,t)$ is Gaussian, and elements of $h(x,t)$ exhibit Gaussian profiles over $x$.

### v4

# Place Cells as Multi-Scale Position Embeddings: Random Walk Transition Kernels for Path Planning

@@ Abstract
The hippocampus supports spatial navigation by encoding cognitive maps through collective place cell activity. We model the place cell population as non-negative spatial embeddings derived from the spectral decomposition of multi-step random walk transition kernels. In this framework, inner product or equivalently Euclidean distance between embeddings encode similarity between locations in terms of their transition probability across multiple scales, forming a cognitive map of adjacency. The combination of non-negativity and inner-product structure naturally induces sparsity, providing a principled explanation for the localized firing fields of place cells without imposing explicit constraints. The temporal parameter that defines the diffusion scale also determines field size, aligning with the hippocampal dorsoventral hierarchy. Our approach constructs global representations efficiently through recursive composition of local transitions, enabling smooth, trap-free navigation and preplay-like trajectory generation. Moreover, theta phase arises intrinsically as the angular relation between embeddings, linking spatial and temporal coding within a single representational geometry.
@@ 1 Introduction
Place cells in the hippocampus are central to spatial cognition, firing selectively at specific locations as animals navigate their environment [1, 2]. Rather than modeling individual place fields, we propose viewing place cell populations as non-negative position embeddings derived from spectral decomposition of multi-step transition kernels that collectively encode spatial relationships. This population-level approach captures the hippocampus’s role in forming cognitive maps, integrating metric and topological properties for flexible navigation [3].
Hippocampal place cells exhibit dynamic behaviors, adapting firing fields to environmental changes [4], scaling along the dorsoventral axis [5], and engaging in preplay before exploration [6]. These properties suggest a multi-scale, adaptive representation that supports navigation across diverse contexts—from narrow passages to open terrains. The hippocampus’s ability to predict novel paths, as in preplay, and maintain robust navigation under environmental deformations further underscores its computational sophistication [7, 8, 9].
We formalize these insights by modeling place cell population as vector embeddings derived from a multi-step symmetric random walk. The transition matrix admits spectral decomposition, enabling construction of non-negative embeddings $h(x,\tau)$ such that the inner product between embeddings approximates the normalized transition probability: $\langle h(x,\tau),h(y,\tau)\rangle=q(y|x,\tau),$ where $h(x,\tau)$ is the embedding at location $x$, and $q(y|x,\tau)$ is the symmetric transition probability over time step $\tau$. Remarkably, combining this inner product structure with non-negativity and orthogonality constraints induces emergent sparsity—the embeddings naturally develop disjoint support sets, explaining why place cells exhibit localized firing fields without explicit regularization. The time parameter $\sqrt{\tau}$ defines a spatial scale hierarchy, mirroring hippocampal dorsoventral scaling [5]. $q(y|x,\tau)$ defines spatial adjacency between $x$ and $y$ at scale or resolution $\sqrt{\tau}$, and the pairwise adjacency relationships $(q(y|x,\tau),\forall x,y)$ are reduced into individual embeddings $(h(x,\tau),\forall x)$ that collectively form a multi-scale, Euclideanized, and sparsified cognitive map of the environment. The Euclideanization emerges from inner product expansion of the transition kernel, the sparsification arises naturally from non-negativity of the place cell responses and the orthogonality between embeddings with zero inner product, and the multi-scale hierarchy enables adaptive navigation across spatial resolutions. Efficient matrix squaring ($P_{2\tau}=P_{\tau}^{2}$) computes global transitions from local ones ($P_{1}$) without past trajectory memorization, enabling preplay-like shortcut detection [10, 7].
Our framework uses gradient ascent on $q(y|x,\tau)=\langle h(x,\tau),h(y,\tau)\rangle$ with adaptive scale selection, choosing the time scale maximizing the gradient for trap-free, smooth trajectories. This produces robust navigation with properties like boundary avoidance, diffraction-like passage guidance, aligning with hippocampal navigation [11, 9, 3]. Continuous interpolation ensures smooth gradient fields, supporting natural paths. Due to the Euclideanization, path planning amounts to gradient descent on Euclidean distance at the selected scale, enabling what can be called “straight forward” path planning.
Beyond spatial coding, we propose a novel theoretical framework where theta phase naturally emerges from the population embedding structure, unifying spatial representation with temporal dynamics and offering a principled account of theta phase precession.
Appendix A provides an extensive review of related work.
Our contributions include: (1) reconceptualizing place cells as non-negative embeddings from spectral decomposition encoding transition probabilities at multiple scales; (2) demonstrating emergent sparsity explaining localized place fields; (3) modeling multi-scale, Euclideanized, and sparsified cognitive maps via time scale parameter $\sqrt{\tau}$ for transition probabilities; (4) employing efficient matrix squaring to build up multi-scale spatial relationships; (5) introducing adaptive gradient ascent for Euclideanized path planning at selected scale; (6) proposing theta phase formulation from population embeddings; and (7) demonstrating biological plausibility through properties like preplay. Bridging connectionist models [12, 13] and cognitive map theories [2, 9], our framework offers a scalable, biologically inspired model of hippocampal spatial navigation.
@@ 2 Method
2.1 Multi-Step Random Walk Transition Kernel
The foundation of our approach is a symmetric random walk on a discrete lattice over the environment (e.g., a $40\times 40$ lattice), with a subset of lattice points belonging to the obstacles. This random walk serves as a mapping policy rather than a goal-reaching policy. Remarkably, this purely random exploration policy leads to optimal path planning without explicit policy optimization. We use $\tau$ to denote the time step of this random walk mapping policy, to avoid confusion with the time $t$ of the planned trajectory. $\tau$ plays the role of scale, which is adaptively selected during path planning.
@@ 2.1 Multi-Step Random Walk Transition Kernel
  • For each unobstructed neighbor $y$ of $x$: $p(y|x,\tau=1)=p_{\text{move}}$ (e.g., 1/4 in the case of 4 nearest neighbors)
  • For self-transition: $p(x|x,\tau=1)=1-|N(x)|\cdot p_{\text{move}}$
@@
where $|N(x)|$ is the number of unobstructed neighbors of $x$. A critical property of this formulation is that the transition probabilities are symmetric, meaning $p(y|x,\tau=1)=p(x|y,\tau=1)$ for all locations $x$ and $y$. This symmetry ensures that the random walk process preserves the bidirectional nature of spatial relationships, which is essential for creating a well-defined proximity metric.
@@
The above random walk defines $\tau$-step symmetric transition kernel $p(y|x,\tau)$ over time step $\tau$, which measures the spatial adjacency between $x$ and $y$ at a spatial scale or resolution captured by $\tau$.
2.2 Heat Diffusion, Geodesic Distance and Topological Connectivity
@@
Our discrete random walk model establishes a connection to the continuous heat equation. For small spatial discretization $dx$ and temporal discretization $dt$ with $dx=\sqrt{dt}$, our discrete random walk converges to the heat diffusion equation with reflecting boundary conditions as $dx\rightarrow 0$ [14, 15]:
@@
$\frac{\partial p(y|x,\tau)}{\partial\tau}=\alpha\nabla^{2}p(y|x,\tau)\quad\text{on}\quad\Omega\setminus\Omega_{\text{obstacles}}$ (1)
$\frac{\partial p(y|x,\tau)}{\partial n}=0\quad\text{on}\quad\partial\Omega_{\text{obstacles}}$ (2)
@@
where $\Omega$ is the whole region, $\Omega_{\text{obstacles}}$ is the region of obstacles, and $\partial\Omega_{\text{obstacles}}$ is its boundary. $\alpha$ is the diffusion coefficient (different one-step transition $p(y|x,\tau=1)$ in discrete case leads to different $\alpha$, e.g., $\alpha=1/4$ for 4-nearest neighbor random walk).
@@
A fundamental result from heat diffusion theory relates the short-time behavior of the heat kernel to geodesic distance. For small values of $\tau$, the heat kernel has the asymptotic form [16, 17]:
$p(y|x,\tau)\approx\frac{1}{4\pi\alpha\tau}\exp\left(-\frac{d_{g}^{2}(x,y)}{4\alpha\tau}\right)$ (3)
@@
where $d_{g}(x,y)$ is the geodesic distance. For open domains, $d_{g}(x,y)$ becomes Euclidean distance, and $p(y|x,\tau)\sim{\cal N}(x,2\alpha\tau)$, a Gaussian distribution with variance $2\alpha\tau$ or standard deviation $\sqrt{2\alpha\tau}$, demonstrating that $\sqrt{\tau}$ correspond to the spatial scale. We take square root of $\tau$ to emphasize this scaling relationship between time and space.
This result demonstrates that $d_{t}(x,y)=-\tau\log q(y|x,\tau)$ approximates the squared geodesic distance for small values of $\tau$, providing a fundamental connection between random walks and geodesic distances in complex environments.
@@
As $\tau$ increases, the transition probability incorporates additional information about path multiplicity and global connectivity, creating a multi-scale representation of spatial relationships. In particular, for large $\tau$, the transition probability is dominated by eigenvalues that are close to 1, whose eigenvectors depend on topological connectedness rather than local geometry.
@@
Appendix D provides detailed background on heat equation. Appendix G explains topological properties for large $\tau$.
2.3 Place Cells as Non-negative Spectral Embeddings
@@
We formalize place cell population as vector embedding $h(x,\tau)\in\mathbb{R}^{n}$, where $n$ is the number of place cells (e.g., $n=500$). The inner product between embeddings approximates the normalized transition probability kernel:
$\langle h(x,\tau),h(y,\tau)\rangle\approx q(y|x,\tau)$ (4)
@@
where $q(y|x,\tau)={p(y|x,\tau)}/{\sqrt{p(x|x,\tau)\cdot p(y|y,\tau)}}$ is the normalized transition probability, so that $\|h(x,\tau)\|=1$ due to normalization. For each cell $i$, $h_{i}(x,\tau)\geq 0$ for biological plausibility, and $h_{i}(x,\tau)$ is the response map or profile of cell $i$ at spatial scale $\sqrt{\tau}$.
@@
The above formulation can be derived through spectral decomposition of the transition matrix. Since the one-step transition matrix $P_{1}$ is symmetric by construction, its powers $P_{\tau}=P_{1}^{\tau}$ admit an eigendecomposition [18]: $P_{\tau}=Q\Lambda^{\tau}Q^{T},$ where $Q$ is orthogonal ($Q^{T}Q=I$) and $\Lambda=\text{diag}(\lambda_{1},\ldots,\lambda_{n})$ contains eigenvalues $0\leq\lambda_{i}\leq 1$ (assuming the random walk is irreducible and aperiodic).
@@
From this spectral decomposition, we can construct a spectral embedding by defining: $H_{i}(x,\tau)=\lambda_{i}^{\tau/2}Q_{i}(x)$ where $Q_{i}$ is the $i$-th column of $Q$. This yields an exact representation of the transition probability through inner products:
$p(y|x,\tau)=\sum_{i}H_{i}(x,\tau)H_{i}(y,\tau)=\langle H(x,\tau),H(y,\tau)\rangle$ (5)
@@
where $H_{i}(x,\tau)$ is the $i$-th element of the vector $H(x,\tau)$. For the normalized transition probability $q(y|x,\tau)$, we define normalized embeddings:
$h_{\text{spec}}(x,\tau)=\frac{H(x,\tau)}{\|H(x,\tau)\|}=\frac{H(x,\tau)}{\sqrt{p(x|x,\tau)}}$ (6)
@@
$\langle h_{\text{spec}}(x,\tau),h_{\text{spec}}(y,\tau)\rangle=\frac{p(y|x,\tau)}{\sqrt{p(x|x,\tau)\cdot p(y|y,\tau)}}=q(y|x,\tau)$ (7)
$p(x|x,\tau)$ is constant in the open field and is smooth in general, so $q$ is essentially a scaled version of $p$.
However, these spectral embeddings may contain negative components, conflicting with the biological constraint that neural firing rates must be non-negative. Horn’s theorem [19, 20], which is built upon the above spectral decomposition, guarantees the existence of a non-negative $h(x,\tau)$, such that $\langle h(x,\tau),h(y,\tau)\rangle=q(y|x,\tau)$ for non-negative matrix factorization.
This formulation represents a fundamental shift from modeling individual place cells to modeling the place cell population as distributed position embedding. The embedding vector $h(x,\tau)$ represents the firing rates of all place cells at location $x$ for spatial scale $\sqrt{\tau}$, capturing the idea that it is the pattern across the population—not the activity of any single cell—that encodes location.
The time parameter $\sqrt{\tau}$ serves as a fundamental unit of spatial resolution or scale, with larger values producing broader, more diffuse representations, and smaller values producing more localized representations. This naturally mirrors the variation in place field sizes observed along the dorsoventral axis of the hippocampus [5, 21].
See Appendix C for more details on spectral analysis. Appendix E provides analytical results for open field, where $q(y|x,\tau)$ is Gaussian, and elements of $h(x,\tau)$ exhibit Gaussian profiles over $x$.